OpenStudy (anonymous):

What is the solution of the equation 4x + 17 = 45 ?

OpenStudy (eric_d):

4x+17=45 4x=45-17 4x=

OpenStudy (eric_d):



OpenStudy (anonymous):

4x=28 which would then equal 7 right?

OpenStudy (eric_d):

yes

OpenStudy (eric_d):

You got it rite